OpenStudy (anonymous):

Solve 2x + 3 > 1 and 5x – 9 ≤ 6.

OpenStudy (paxpolaris):

solve each inequality separately, then combine

OpenStudy (paxpolaris):

\[\large 2x + 3 > 1\]\[\implies \large 2x > 1-3\]\[\implies \large 2x > -2\] \[\implies \LARGE x > -1\]\[\implies \LARGE -1<x\]

OpenStudy (paxpolaris):

what do you get after solving: \(\Large 5x – 9 \le 6\) ??

OpenStudy (anonymous):

2x + 3 > 1 2x + 3 - 3 > 1 - 3 2x > -2 x > -1 && 5x - 9 <= 6 5x - 9 + 9 <= 6 + 9 5x <= 15 x <= 3 so answer is -1 < x <= 3